When a strong acid-strong base neutralization reaction is also a limiting reactant situation, the leftover reactant will determine your final solution pH. Remember that a limiting reactant indicates that you will run out of one reactant first.

Considering the reaction of potassium hydroxide and hydrobromic acid, notice that if you have excess acid leftover it would react with water to produce H3O+ ions in solution: HBr + H2O → Br- + H3O+

If you have excess base left over, it will dissociate in water to produce OH- ions: KOH → K+ + OH-

You mix 2 moles of HBr with 3 moles of KOH in enough water to make 1 L of solution. What is the pH of your final solution?

I'm really struggling on how to find the limiting reactant and find pH from there
Answer #1

As HBr has less moles so it is limiting reagent.

KOH + HBr----> KBr + H2O

So excess KOH = 3-2 = 1.0 mol

[OH-] = 1/1 = 1 M

pOH = -log(H+]

= 0

pH = 14